At approximately 08:00 this morning mains power to our primary data centre was lost. Service was not disrupted as the UPS and Diesel generators provided continuity of power. However, at approximately 14:46 this afternoon a fault occurred that result in a loss of power from the diesel generators. Onsite engineers… Continue Reading Power Outage 2017/01/21
Earlier today we began to see reports of issues contacting outlook.com, hotmail.com and associated addresses for the free mail services offered by Microsoft. After investigations we could see this was only impacting a subset of mail flow to these domains. Further to this the failures were of such an intermittent… Continue Reading [INFO] Hotmail/Outlook email delivery – ‘540 Transaction failed’
There have been recurring performance issues encountered today with the webservices on vlinweb4. Engineers are currently working to isolate the source of the issue and restore normal service.
We encountered a corrupt configuration within one of the nodes servicing ‘relay2’ which resulted in some emails being rejected from authorised sources. The underlying issue was resolved yesterday evening and changes have been put in place to the maintenance process that are designed to prevent it from reoccurring.
We have received notification that there may be a disruption of 1 hour between 00:00 and 04:00 tomorrow (2015-01-15) due to a router upgrade. This will impact a subset of all web, email and VM hosting services.
Our Second DC provider is currently encountering network connectivity issues which is impacting services served from the 22.214.171.124 and 126.96.36.199 networks. They are currently investigating.
Routine monitoring has highlighted intermittent access issues for customers with Windows based hosting account. We encountered an issue during the past hour where there was intermittent access to email hosting for customers with Windows based hosting. We have resolved the stability issue and normal access is now restored. We will… Continue Reading Email stability (resolved)
We are currently experiencing an unplanned outage due to power issues at the data centre. We are liaising with our data centre provider and will give an update as soon as we have more information.
A number of services encountered a short period of downtime this evening due to a SAN controller freezing. We have restarted the controller and service has now been restored. We will now begin a detailed investigation to understand the issue fully and prevent the issue from occurring again.
The Cloud Backup service will be unavailable between 00:30 and 12:30 (BST / GMT+1) on Thursday 1st August 2013 to enable a scheduled software update to be released. During this Window it will not be possible to backup or restore data as the update is applied. The new version will… Continue Reading Cloud Backup maintenance reminder